OpenStudy (anonymous):

The result when a number is divided by 2 is equal to the result when that same number is divided by 4. What is that number? Please explain? I know the answer is 0 but I'm not understanding HOW the answer is 0.

OpenStudy (australopithecus):

set up the equality x/2 = x/4

OpenStudy (australopithecus):

and solve for it you will see the answer is zero

OpenStudy (australopithecus):

Also it is pretty easy to see the only number that makes this equality actually equal is 0

OpenStudy (rational):

x/2 = x/4 x/2 - x/4 = 0 2x - x = 0 x = 0

OpenStudy (rational):

its easy to rush to cancel x the moment you see it both sides (from personal experience :P ) but you should NEVER cancel a variable both sides without fully knowing that it never equals 0

OpenStudy (anonymous):

I think I understand...thank you both.